OpenStudy (anonymous):

Solve by the elimination method 7r-9s=45 9r-7s=95

OpenStudy (saifoo.khan):

x = 135/8 y = 65/8

OpenStudy (anonymous):

Saif is that simplified?

OpenStudy (saifoo.khan):

Yup!

OpenStudy (anonymous):

7r-9s=45----------*9 9r-7s=95----------*7 63r-81s=405 63r-49s=665 -32s=-260 s=65/8 9r-7(65/8)=95 9r-455/8=95 9r-56.875=95 r=135/8